# matrix,arrows,decorations.pathmorphing

### v2

@@ 2.3. The case $S=S_{0,5}$
$1\to\langle{\upsilon}\rangle\to\pi_{1}({\mathcal{M}}_{1,[2]})\to\pi_{1}({\mathcal{M}}_{0,1[4]})\to 1$
In terms of mapping class groups, this can be described as follows. Let ${\upsilon}\in{\Gamma}(S_{1,2})$ be the hyperelliptic involution, let $S_{/{\upsilon}}$ be the quotient of the surface $S_{1,2}$ by ${\upsilon}$ and let $B_{\upsilon}$ be the branch locus of the orbit map $S\to S_{/{\upsilon}}$. The surface $S_{/{\upsilon}}$ is a $1$-punctured sphere, there is a diffeomorphism $S_{/{\upsilon}}\smallsetminus B_{\upsilon}\cong S_{0,5}$ and, if we denote by $Q$ the puncture of $S_{0,5}$ which corresponds to the puncture of $S_{/{\upsilon}}$ via the above diffeomorphism, by Birman-Hilden theory, there is a short exact sequence (cf. Theorem 2.3 in [8]):
$1\to\langle{\upsilon}\rangle\to{\Gamma}(S_{1,2})\to{\Gamma}(S_{0,5})_{Q}\to 1,$
where ${\Gamma}(S_{0,5})_{Q}$ is the stabilizer of the puncture $Q$ in ${\Gamma}(S_{0,5})$. There is then a natural isomorphism between this short exact sequence and the one obtained above in a geometric way. In particular, there is a natural isomorphism $\mathrm{P}\Gamma(S_{1,2})\cong{\Gamma}(S_{0,5})_{Q}$ so that $\mathrm{P}\Gamma(S_{0,5})$ identifies with a subgroup of $\mathrm{P}\Gamma(S_{1,2})$. We record the following for future use:
@@
The image of a nonseparating Dehn twist via the epimorphism ${\Gamma}(S_{1,2})\to{\Gamma}(S_{0,5})_{Q}$ is a braid twist. Hence, the image of the square of a nonseparating Dehn twist is a Dehn twist about a simple closed curve on $S_{0,5}$ which bounds a $2$-punctured disc. The subgroup of ${\Gamma}(S_{1,2})$ generated by squares of nonseparating Dehn twists is contained in $\mathrm{P}\Gamma(S_{1,2})$ and has trivial intersection with $\langle{\upsilon}\rangle$. Hence, it identifies with the subgroup of ${\Gamma}(S_{0,5})_{Q}$ generated by Dehn twists about simple closed curves on $S_{0,5}$ bounding $2$-punctured discs, which is $\mathrm{P}\Gamma(S_{0,5})$. ∎
@@
From the above discussion, we see, in particular, that there is a natural finite étale morphism ${\mathcal{M}}_{0,5}\to{\mathcal{M}}_{1,2}$, so that ${\mathcal{M}}_{0,5}$ identifies with a level structure over ${\mathcal{M}}_{1,[2]}$. Moreover, the DM compactification ¯${\overline{\mathcal{M}}}_{0,5}$ of ${\mathcal{M}}_{0,5}$ coincides with the DM compactification of the latter as a level structure. Therefore, there are natural isomorphisms ${\mathbb{M}}(S_{0,5})\cong{\mathbb{M}}(S_{1,2})$, $\overline{{\mathbb{M}}}(S_{0,5})\cong\overline{{\mathbb{M}}}(S_{1,2})$ and then $\partial\overline{{\mathbb{M}}}(S_{0,5})\cong\partial\overline{{\mathbb{M}}}(S_{1,2})$.
@@
The Teichmüller spaces ${\mathcal{T}}(S_{0,5})$ and ${\mathcal{T}}(S_{1,2})$ are the universal covers of ${\mathcal{M}}_{0,5}$ and ${\mathcal{M}}_{1,2}$, respectively. Hence, there is a natural ${\Gamma}(S_{1,2})$-equivariant isomorphism ${\mathcal{T}}(S_{0,5})\cong{\mathcal{T}}(S_{1,2})$. There is also a compatible ${\Gamma}(S_{1,2})$-equivariant isomorphism of curve complexes $C(S_{0,5})\cong C(S_{1,2})$. Since the Weil-Petersson metric on the Teichmüller space is determined, modulo strong equivalence, by the Fenchel-Nielsen coordinates, it follows that the isomorphism ${\mathcal{T}}(S_{0,5})\cong{\mathcal{T}}(S_{1,2})$ induces a natural ${\Gamma}(S_{1,2})$-equivariant isomorphism of augmented Teichmüller spaces $\overline{{\mathcal{T}}}(S_{0,5})\cong\overline{{\mathcal{T}}}(S_{1,2})$ and then $\partial\overline{{\mathcal{T}}}(S_{0,5})\cong\partial\overline{{\mathcal{T}}}(S_{1,2})$.
@@
Even though the pants graph $C_{P}(S_{1,2})$ is the $1$-skeleton of a ${\Gamma}(S_{1,2})$-equivariant triangulation of $\partial\overline{{\mathcal{T}}}(S_{1,2})$, the same holds for $C_{P}(S_{0,5})$ and $\partial\overline{{\mathcal{T}}}(S_{0,5})$ and the vertex sets of the pants graphs $C_{P}(S_{1,2})$ and $C_{P}(S_{0,5})$ naturally identify, there is no natural map between these two graphs to account for it. This follows from a general result by Aramayona (cf. Theorem A in [1]) but also from a careful analysis of the pairs of curves with minimal intersection occurring in $S_{1,2}$ and $S_{0,5}$.
@@
In any case, from the natural ${\Gamma}(S_{1,2})$-equivariant bijective map of 00-simplices $C_{P}(S_{1,2})_{0}\cong C_{P}(S_{0,5})_{0}$, passing to ${\check{\Gamma}}(S_{1,2})$-completions, we get a natural ${\check{\Gamma}}(S_{1,2})$-equivariant bijective map ${\check{C}}_{P}(S_{1,2})_{0}\cong{\check{C}}_{P}(S_{0,5})_{0}$. By Proposition 8.3 in [6], there is a natural monomorphism $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{1,2}))$ and then $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$. From Lemma 2.4, it then follows that every continuous automorphism of ${\check{C}}_{P}(S_{1,2})$ induces one of ${\check{C}}_{P}(S_{0,5})$ (compatible on vertex sets with the ${\check{\Gamma}}(S_{1,2})$-equivariant bijection given above).
@@
Let $\{{\sigma}_{0},{\sigma}_{1}\}\subset C(S)_{d(S)-1}\subset{\check{C}}(S)_{d(S)-1}$ be the vertex set of $e$. By the first item of Lemma 6.6 in [6], the edge $e$ is then contained in the profinite Farey subgraph ${\widehat{\operatorname{F}}}_{{\sigma}_{0}\cap{\sigma}_{1}}$, which is obtained as the $\widehat{\operatorname{PSL}_{2}(\mathbb{Z})}$-completion of the Farey subgraph $\rm F_{{\sigma}_{0}\cap{\sigma}_{1}}\subset C_{P}(S)$. Hence it is enough to prove the statement of the lemma for the profinite Farey graph ${\widehat{\operatorname{F}}}$.
Given two vertices $v_{0},v_{1}\in\rm F_{0}$, there is a unique geodesic ${\gamma}$ in $\overline{{\mathbb{H}}}$ connecting these two points and a finite index subgroup ${\Gamma}^{\lambda}$ of $\operatorname{PSL}_{2}(\mathbb{Z})$ such that, for all ${\Gamma}^{{\lambda}^{\prime}}\leq{\Gamma}^{\lambda}$, the image of ${\gamma}$ in the quotient surface ¯$\overline{{\mathbb{H}}}/{\Gamma}^{{\lambda}^{\prime}}$ is a simple geodesic arc. This implies that, if $\{v_{0},v_{1}\}$ is the vertex set of an edge of ${\widehat{\operatorname{F}}}$, the distance between $v_{0}$ and $v_{1}$ in $\overline{{\mathbb{H}}}$ is $1$, which is possible only if $\{v_{0},v_{1}\}$ is the vertex set of an edge of $\rm F$. ∎
From Lemma 2.6, it follows that $f$ induces an automorphism of the pants complex $C_{P}(S_{1,2})$. By Theorem 1 and Theorem 2 of [19], we then have that $f\in\operatorname{Inn}({\Gamma}^{\pm}(S_{1,2}))$ which completes the proof of the case $S=S_{1,2}$ of Proposition 2.1.
2.5. Four lemmas
The following definition will play a fundamental role in the proof, by induction, of the general case of Proposition 2.1:
@@ 2.5. Four lemmas
For $d(S)>1$, every $(d(S)-2)$-multicurve on $S$ contains at least a simple closed curve which is either nonseparating or bounds a $2$-punctured disc. For a fixed such simple closed curve ${\gamma}$, we then let $L_{\gamma}$ be the closed subgraph of ${\check{C}}_{P}(S)$ which is the union of all profinite Farey subgraphs ${\widehat{\operatorname{F}}}_{\sigma}$ such that ${\gamma}\in{\sigma}$.
@@
By Remark 4.7 in [4], the link $\operatorname{Lk}({\gamma})\subset{\check{C}}(S)$ is naturally isomorphic to ${\check{C}}(S_{\gamma})$. Let $\xi\colon\thinspace{\check{C}}(S_{\gamma})\stackrel{{\scriptstyle\sim}}{{\to}}\operatorname{Lk}({\gamma})$ be such isomorphism. We can then identify the vertex set of ${\check{C}}_{P}(S_{\gamma})$ with a subset of the vertex set of ${\check{C}}_{P}(S)$ by sending a $(d(S_{\gamma})-1)$-simplex ${\sigma}$ of ${\check{C}}(S_{\gamma})$ to the $(d(S)-1)$-simplex $\xi({\sigma})\cup\{{\gamma}\}$ of $\operatorname{Star}({\gamma})\subset{\check{C}}(S)$. The image of this map is precisely the vertex set of the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S)$ and it is easy to check that it induces a map between the edges of ${\check{C}}_{P}(S_{\gamma})$ and those of $L_{\gamma}$, from which the conclusion follows. ∎
The procongruence curve complex ${\check{C}}(S)$ can be identified with the simplicial profinite complex whose simplices are the inertia groups $\{{\hat{\mathrm{I}}}_{\sigma}\}_{{\sigma}\in{\check{C}}(S)}\subset{\check{\Gamma}}(S)$ (cf. Remark 4.13 in [6]), so that there is a natural continuous action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on ${\check{C}}(S)$:
@@
By Theorem 7.3 in [6], the kernel of the homomorphism $\check{\Theta}_{\mathbb{I}}$ is contained in the subgroup $\operatorname{Hom}(\mathrm{P}{\check{\Gamma}}(S)/Z(\mathrm{P}{\check{\Gamma}}(S)),Z(\mathrm{P}{\check{\Gamma}}(S)))$ of $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$ described in Lemma 7.4 and Lemma 3.5 in [6], where we denote by $Z(\mathrm{P}{\check{\Gamma}}(S))$ the center of $\mathrm{P}{\check{\Gamma}}(S)$. From its explicit description, it follows that the intersection of $\operatorname{Hom}(\mathrm{P}{\check{\Gamma}}(S)/Z(\mathrm{P}{\check{\Gamma}}(S)),Z(\mathrm{P}{\check{\Gamma}}(S)))$ with the subgroup $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ inside $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$ is trivial, which implies the lemma. ∎
@@
An automorphism of $\mathrm{P}{\check{\Gamma}}(S)$ which preserves the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ also preserves its centralizer $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})$ in $\mathrm{P}{\check{\Gamma}}(S)$ and, since, for ${\sigma}\in\operatorname{Lk}({\gamma})$, the inertia group ${\hat{\mathrm{I}}}_{\sigma}$ is contained in $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})$, the action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}$ on $\operatorname{Lk}({\gamma})$ factors through the homomorphism induced by restriction:
@@
By Corollary 4.12 in [6], there is a natural isomorphism $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})\cong\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$, so that we can identify $\operatorname{Aut}^{\mathbb{I}}(Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma}))$ with the closed subgroup $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}})$ of $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}})$ consisting of those automorphisms which preserve the set of conjugacy classes of the procyclic subgroups of $\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$ generated by profinite Dehn twists.
By Theorem 4.10 in [6], there are the exact sequences:
[…]
where the homomorphism $\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}\to\{\pm 1\}$ is induced by the action of the stabilizer $\mathrm{P}\Gamma(S)_{\gamma}$ on the orientation of the simple closed curve ${\gamma}$.
@@
By Theorem 4.14 in [6], the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ is the center of $\mathrm{P}{\check{\Gamma}}(S)_{\vec{\gamma}}$, hence, it is preserved by every element of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{\vec{\gamma}})$ and there is a natural homomorphism:
@@
The following simple lemma in group theory will also be useful:

### v3

@@ 2.2. Automorphisms of the procongruence pants complex and orientations
(6) $1\to\mathbb{Z}/2\to\pi_{1}({\mathcal{M}}_{1,[2]})\to\pi_{1}({\mathcal{M}}_{0,1[4]})\to 1,$
which, in terms of mapping class groups, can be described as follows.
Let ${\upsilon}\in{\Gamma}(S_{1,2})$ be the hyperelliptic involution, let $S_{/{\upsilon}}$ be the quotient of the surface $S_{1,2}$ by $\upsilon$} and let $B_{\upsilon}$ be the branch locus of the orbit map $S\to S_{/{\upsilon}}$. The surface $S_{/{\upsilon}}$ is a $1$-punctured sphere, there is a diffeomorphism $S_{/{\upsilon}}\smallsetminus B_{\upsilon}\cong S_{0,5}$ and, if we denote by $Q$ the puncture of $S_{0,5}$ which corresponds to the puncture of $S_{/{\upsilon}}$ via the above diffeomorphism, by Birman–Hilden theory, there is a split short exact sequence (cf. [11, Theorem 2.3], for instance):
(7) $1\to\langle{\upsilon}\rangle\to{\Gamma}(S_{1,2})\to{\Gamma}(S_{0,5})_{Q}\to 1,$
where ${\Gamma}(S_{0,5})_{Q}$ is the stabilizer of the puncture $Q$ in ${\Gamma}(S_{0,5})$ and the splitting is given by internal direct product decomposition ${\Gamma}(S_{1,2})=\langle{\upsilon}\rangle\cdot\mathrm{P}\Gamma(S_{1,2})$. There is then a natural isomorphism between the split short exact sequences (6) and (7).
In particular, the epimorphism ${\Gamma}(S_{1,2})\to{\Gamma}(S_{0,5})_{Q}$ induces an isomorphism $\mathrm{P}\Gamma(S_{1,2})\cong{\Gamma}(S_{0,5})_{Q}$ which identifies $\mathrm{P}\Gamma(S_{0,5})$ with a subgroup of $\mathrm{P}\Gamma(S_{1,2})$. We record the following for future use:
@@
The image of a nonseparating Dehn twist via the epimorphism ${\Gamma}(S_{1,2})\to{\Gamma}(S_{0,5})_{Q}$ is a braid twist. Hence, the image of the square of a nonseparating Dehn twist is a Dehn twist about a simple closed curve on $S_{0,5}$ which bounds a $2$-punctured disc. The subgroup of ${\Gamma}(S_{1,2})$ generated by squares of nonseparating Dehn twists is contained in $\mathrm{P}\Gamma(S_{1,2})$ and has trivial intersection with $\langle{\upsilon}\rangle$. Therefore, it identifies with the subgroup of ${\Gamma}(S_{0,5})_{Q}$ generated by Dehn twists about simple closed curves on $S_{0,5}$ bounding $2$-punctured discs, which is $\mathrm{P}\Gamma(S_{0,5})$. ∎
@@
From the isomorphism ${\mathcal{M}}_{1,2}\cong{\mathcal{M}}_{0,1[4]}$, it also follows that there is a natural finite étale morphism ${\mathcal{M}}_{0,5}\to{\mathcal{M}}_{1,2}$, so that ${\mathcal{M}}_{0,5}$ identifies with a level structure over ${\mathcal{M}}_{1,[2]}$. Moreover, the DM compactification ${\overline{\mathcal{M}}}_{0,5}$ of ${\mathcal{M}}_{0,5}$ coincides with the DM compactification of the latter as a level structure. Therefore, there are natural isomorphisms ${\mathbb{M}}(S_{0,5})\cong{\mathbb{M}}(S_{1,2})$, $\overline{{\mathbb{M}}}(S_{0,5})\cong\overline{{\mathbb{M}}}(S_{1,2})$ and then $\partial\overline{{\mathbb{M}}}(S_{0,5})\cong\partial\overline{{\mathbb{M}}}(S_{1,2})$.
@@
The Teichmüller spaces ${\mathcal{T}}(S_{0,5})$ and ${\mathcal{T}}(S_{1,2})$ are the universal covers of ${\mathcal{M}}_{0,5}$ and ${\mathcal{M}}_{1,2}$, respectively. Hence, there is a natural ${\Gamma}(S_{1,2})$-equivariant isomorphism ${\mathcal{T}}(S_{0,5})\cong{\mathcal{T}}(S_{1,2})$. There is also a compatible ${\Gamma}(S_{1,2})$-equivariant isomorphism of curve complexes $C(S_{0,5})\cong C(S_{1,2})$. Since the Weil–Petersson metric on the Teichmüller space is determined, modulo strong equivalence, by the Fenchel–Nielsen coordinates, it follows that the isomorphism ${\mathcal{T}}(S_{0,5})\cong{\mathcal{T}}(S_{1,2})$ induces a natural ${\Gamma}(S_{1,2})$-equivariant isomorphism of augmented Teichmüller spaces $\overline{{\mathcal{T}}}(S_{0,5})\cong\overline{{\mathcal{T}}}(S_{1,2})$ and then $\partial\overline{{\mathcal{T}}}(S_{0,5})\cong\partial\overline{{\mathcal{T}}}(S_{1,2})$.
@@
Even though the pants graph $C_{P}(S_{1,2})$ is the $1$-skeleton of a ${\Gamma}(S_{1,2})$-equivariant triangulation of $\partial\overline{{\mathcal{T}}}(S_{1,2})$, the same holds for $C_{P}(S_{0,5})$ and $\partial\overline{{\mathcal{T}}}(S_{0,5})$, and the vertex sets of the pants graphs $C_{P}(S_{1,2})$ and $C_{P}(S_{0,5})$ naturally identify, there is no natural map between these two graphs to account for it. This follows from a general result by Aramayona (cf. [1, Theorem A]) but also from a careful analysis of the pairs of curves with minimal intersection occurring in $S_{1,2}$ and $S_{0,5}$.
@@
In any case, from the natural ${\Gamma}(S_{1,2})$-equivariant bijective map of 0-simplices $C_{P}(S_{1,2})_{0}\cong C_{P}(S_{0,5})_{0}$, passing to ${\check{\Gamma}}(S_{1,2})$-completions, we get a natural ${\check{\Gamma}}(S_{1,2})$-equivariant bijective map ${\check{C}}_{P}(S_{1,2})_{0}\cong{\check{C}}_{P}(S_{0,5})_{0}$. By [9, Proposition 8.3], there is a natural monomorphism $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{1,2}))$ and so $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$. Lemma 2.5 then implies that every continuous automorphism of ${\check{C}}_{P}(S_{1,2})$ induces one of ${\check{C}}_{P}(S_{0,5})$ (compatible on vertex sets with the ${\check{\Gamma}}(S_{1,2})$-equivariant bijection given above).
@@
Let $\{{\sigma}_{0},{\sigma}_{1}\}\subset C(S)_{d(S)-1}\subset{\check{C}}(S)_{d(S)-1}$ be the vertex set of $e$. By the first item of [9, Lemma 6.6], the edge $e$ is then contained in the profinite Farey subgraph ${\widehat{\operatorname{F}}}_{{\sigma}_{0}\cap{\sigma}_{1}}$, which is obtained as the $\widehat{\operatorname{PSL}_{2}(\mathbb{Z})}-$completion of the Farey subgraph ${\rm F}_{{\sigma}_{0}\cap{\sigma}_{1}}\subset C_{P}(S)$. Hence it is enough to prove the statement of the lemma for the profinite Farey graph $^{\widehat{\operatorname{F}}}$.
Given two vertices $v_{0},v_{1}\in\rm F_{0}$, by basic plane hyperbolic geometry, we know that there is a unique geodesic $\gamma$} in $\overline{{\mathbb{H}}}$ connecting these two points and a finite index subgroup ${\Gamma}^{\lambda}$ of $\operatorname{PSL}_{2}(\mathbb{Z})$ such that, for all ${\Gamma}^{{\lambda}^{\prime}}\leq{\Gamma}^{\lambda}$, the image of $\gamma$} in the quotient surface $\overline{{\mathbb{H}}}/{\Gamma}^{{\lambda}^{\prime}}$ is a simple geodesic arc. This implies that, if $\{v_{0},v_{1}\}$ is the vertex set of an edge of $^{\widehat{\operatorname{F}}}$, the distance between $v_{0}$ and $v_{1}$ in $\overline{{\mathbb{H}}}$ is $1$, which is possible only if $\{v_{0},v_{1}\}$ is the vertex set of an edge of $\rm F$. ∎
From Lemma 2.7, it follows that $f$ induces an automorphism of the pants complex $C_{P}(S_{1,2})$. By [29, Theorem 1 and Theorem 2], we then have that $f\in\operatorname{Inn}({\Gamma}^{\pm}(S_{1,2}))$ which completes the proof of the case $S=S_{1,2}$ of Theorem 2.2.
2.5. Three lemmas
The following definition will play a fundamental role in the proof, by induction, of the general case of Theorem 2.2:
@@ 2.5. Three lemmas
For $d(S)>1$, every (d(S)2)(d(S)-2)-multicurve on $S$ contains at least a simple closed curve which is either nonseparating or bounds a $2$-punctured disc. For a fixed such simple closed curve $\gamma$}, we then let $L_{\gamma}$ be the closed subgraph of ${\check{C}}_{P}(S)$ which is the union of all profinite Farey subgraphs ${\widehat{\operatorname{F}}}_{\sigma}$ such that ${\gamma}\in{\sigma}$.
@@
By [6, Remark 4.7], the link $\operatorname{Lk}({\gamma})\subset{\check{C}}(S)$ is naturally isomorphic to ${\check{C}}(S_{\gamma})$. Let $\xi\colon\thinspace{\check{C}}(S_{\gamma})\stackrel{{\scriptstyle\sim}}{{\to}}\operatorname{Lk}({\gamma})$ be such an isomorphism. We can then identify the vertex set of ${\check{C}}_{P}(S_{\gamma})$ with a subset of the vertex set of ${\check{C}}_{P}(S)$ by sending a $(d(S_{\gamma})-1)$-simplex $\sigma$} of ${\check{C}}(S_{\gamma})$ to the (d(S)1)(d(S)-1)-simplex $\xi({\sigma})\cup\{{\gamma}\}$ of $\operatorname{Star}({\gamma})\subset{\check{C}}(S)$. The image of this map is precisely the vertex set of the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S)$. In order to prove that it also induces a map between the edges of ${\check{C}}_{P}(S_{\gamma})$ and those of $L_{\gamma}$, let us observe the following. Let $\mathring{L}_{\gamma}\subset C_{P}(S)$ be the union of all Farey subgraphs $\operatorname{F}_{\sigma}$ such that ${\gamma}\in{\sigma}$. Then, $\mathring{L}_{\gamma}$ is a dense subgraph of $L_{\gamma}$ and it is clear that the map defined above maps the edge set of the pants complex $C_{P}(S_{\gamma})$ onto the edge set of $\mathring{L}_{\gamma}$. Since the map defined above is ${\check{\Gamma}}(S_{\gamma})$-equivariant, this implies that it maps the edge set of ${\check{C}}_{P}(S_{\gamma})$ onto the edge set of $L_{\gamma}$. ∎
Remark 2.10.
Let ${\mathcal{G}}({\widehat{\Gamma}}(S))$ be the profinite set of closed subgroups of ${\check{\Gamma}}(S)$. By [6, Theorem 6.9], for every $0\leq i\leq d(S)-1$, the assignment ${\sigma}\mapsto{\hat{\mathrm{I}}}_{\sigma}$, for ${\sigma}\in{\check{C}}(S)_{i}$, defines a ${\check{\Gamma}}(S)$-equivariant continuous embedding:
${\mathcal{I}}_{i}\colon\thinspace{\check{C}}(S)_{i}\hookrightarrow{\mathcal{G}}({\check{\Gamma}}(S)).$
The procongruence curve complex ${\check{C}}(S)$ then identifies with the abstract simplicial profinite complex ${\check{C}}_{\mathcal{I}}(S)$ whose set of $h$-simplices is the set of of closed subgroups $\{{\hat{\mathrm{I}}}_{\sigma}\}_{{\sigma}\in{\check{C}}_{h}(S)}$.
By Remark 2.10, there is a natural continuous action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on ${\check{C}}(S)$. We have:
@@
By [9, Theorem 7.3], the kernel of the homomorphism $\check{\Theta}_{\mathbb{I}}$ is contained in the image of the monomorphism $\exp\colon\thinspace\operatorname{Hom}(\mathrm{P}{\check{\Gamma}}(S)/Z(\mathrm{P}{\check{\Gamma}}(S)),Z(\mathrm{P}{\check{\Gamma}}(S)))\to\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$, where we denote by $Z(\mathrm{P}{\check{\Gamma}}(S))$ the center of $\mathrm{P}{\check{\Gamma}}(S)$. This implies the lemma for $S\neq S_{1,1},S_{2}$, since in this case $Z(\mathrm{P}{\check{\Gamma}}(S))=\{1\}$. Otherwise, the center $Z(\mathrm{P}{\check{\Gamma}}(S))$ is generated by the hyperelliptic involution $\upsilon$} and we have $\operatorname{Hom}(\mathrm{P}{\check{\Gamma}}(S)/\langle{\upsilon}\rangle,\langle{\upsilon}\rangle)\cong\operatorname{Hom}(\mathrm{P}\Gamma(S)/\langle{\upsilon}\rangle,\langle{\upsilon}\rangle)\cong\mathbb{Z}/2$.
The automorphism $\exp\phi\in\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$ in the image of $0\neq\phi\in\operatorname{Hom}(\mathrm{P}{\check{\Gamma}}(S)/\langle{\upsilon}\rangle,\langle{\upsilon}\rangle)$ (cf. [9, Lemma 7.4]) maps a nonseparating Dehn twist $\tau_{\gamma}\in\mathrm{P}\Gamma(S)$ to ${\upsilon}\cdot\tau_{\gamma}$. Therefore, $\exp\phi\notin\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$, which implies the lemma. ∎
@@
An automorphism of $\mathrm{P}{\check{\Gamma}}(S)$ which preserves the procyclic inertia group ${\hat{\mathrm{I}}}_{\gamma}$ also preserves its centralizer $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})$ in $\mathrm{P}{\check{\Gamma}}(S)$ and, since, for ${\sigma}\in\operatorname{Lk}({\gamma})$, the inertia group ${\hat{\mathrm{I}}}_{\sigma}$ is contained in $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})$, the action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}$ on $\operatorname{Lk}({\gamma})$ factors through the homomorphism induced by restriction:
@@
By [6, Corollary 6.1], there is a natural isomorphism $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})\cong\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$, so that we can identify $\operatorname{Aut}^{\mathbb{I}}(Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma}))$ with the closed subgroup $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}})$ of $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}})$ consisting of those automorphisms which preserve the set of conjugacy classes of the procyclic subgroups of $\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$ generated by profinite Dehn twists.
Since $\gamma$} is either nonseparating or bounding a 2-punctured disc and the pure mapping class group of $3$-punctured sphere is trivial, by [9, Theorem 4.10], we have the exact sequences:
$1\rightarrow\mathrm{P}{\check{\Gamma}}(S)_{\vec{{\gamma}}}\to\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}\to\{\pm 1\}\hskip 14.22636pt\mbox{ and }\hskip 14.22636pt1\rightarrow\hat{\mathrm{I}}_{\gamma}\to\mathrm{P}{\check{\Gamma}}(S)_{\vec{{\gamma}}}\to\mathrm{P}{\check{\Gamma}}(S_{\gamma})\to 1,$
where the homomorphism $\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}\to\{\pm 1\}$ is induced by the action of the stabilizer $\mathrm{P}\Gamma(S)_{\gamma}$ on the orientation of the simple closed curve $\gamma$}. In particular, for $\gamma$} bounding a $2$-punctured disc, we have $\mathrm{P}{\check{\Gamma}}(S)_{\vec{{\gamma}}}=\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$.
@@
By [9, Theorem 4.14], the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ is the center of $\mathrm{P}{\check{\Gamma}}(S)_{\vec{\gamma}}$, hence, it is preserved by every element of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{\vec{\gamma}})$ and there is a natural homomorphism:
@@
2.6. Connectedness of various curve complexes
For the proof of Theorem 2.2, we need the connectivity of some curve complexes. Even though these results are probably well known to experts, we include a proof for lack of suitable references.
Definition 2.13.
  1. (i)
    For $n\geq 4$, let $C_{b}(S_{0,n})$ be the curve complex defined as the full subcomplex of the complex of curves $C(S_{0,n})$ whose vertices consist of isotopy classes of simple closed curves on $S_{0,n}$ which bound a $2$-punctured disc.
  2. (ii)
    For $n\geq 2$, let $C_{0b}(S_{1,n})$ be the full subcomplex of the curve complex $C(S_{1,n})$ whose vertices are isotopy classes of either nonseparating simple closed curves or simple closed curves which bound a $2$-punctured disc on $S_{0,n}$.